OpenStudy (ragingsquirrel):

What is the reciprocal of 2 3/4?

OpenStudy (anonymous):

2 3/4 = 2 + 3/4 = 11/4 Reciprocal = 4/11

OpenStudy (ragingsquirrel):

Ohhh I think i forgot that but I do remember learning that. Thanks